Skip to main contentSkip to solution

The probability of a man hitting a target is 1/2. How many times must he fire so that the probability of hitting the target at least once is more than 90%?

Solution

✅ Correct Option: 2

The probability of hitting the target is 12\frac{1}{2} and the probability of missing is 12\frac{1}{2}.

The probability of hitting at least once equals 1 minus the probability of missing all shots:

P(hitting at least once)=1−P(missing all shots)P(\text{hitting at least once}) = 1 - P(\text{missing all shots})


If the man fires nn times:

P(missing all n shots)=(12)nP(\text{missing all } n \text{ shots}) = \left(\frac{1}{2}\right)^n

Therefore:

P(hitting at least once)=1−(12)nP(\text{hitting at least once}) = 1 - \left(\frac{1}{2}\right)^n


The probability of hitting at least once must be more than 90%:

1−(12)n>0.901 - \left(\frac{1}{2}\right)^n > 0.90

(12)n<0.10\left(\frac{1}{2}\right)^n < 0.10


For n=3n = 3:

(12)3=18=0.125\left(\frac{1}{2}\right)^3 = \frac{1}{8} = 0.125

P(hitting at least once)=1−0.125=0.875=87.5%P(\text{hitting at least once}) = 1 - 0.125 = 0.875 = 87.5\%

This is less than 90%.


For n=4n = 4:

(12)4=116=0.0625\left(\frac{1}{2}\right)^4 = \frac{1}{16} = 0.0625

P(hitting at least once)=1−0.0625=0.9375=93.75%P(\text{hitting at least once}) = 1 - 0.0625 = 0.9375 = 93.75\%

This is more than 90%.


The man must fire 4 times to ensure the probability of hitting the target at least once is more than 90%.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question